OpenStudy (anonymous):

help me understand these 2 questions are driving me nuts!! >w< 3. Sandra is riding the ferris wheel, and her height can be modeled by the equation h(t)=25cos(pi/14t)+31, where H represents the height of the person above the ground in feet at T seconds. Assuming Sandra begins the ride at the top, how far from the ground is the edge of the Ferris wheel, when Sandra's height above ground reaches a minimum?

OpenStudy (anonymous):

4. Prove : sinθ-sinθ*cos^2θ=sin^3θ

OpenStudy (johnweldon1993):

Interesting... so \[\large h(t) = 25\cos(\frac{\pi}{14}t) + 31\] Since we want the height at a MINIMUM, to me that screams take the derivative and set it equal to 0 So what is the derivative of h(t) ?

OpenStudy (perl):

*

OpenStudy (anonymous):

@perl @johnweldon1993 This is what I got so for number 3 h(t)=25cos(pi/14t)+31 h(t)=25cos+31 h(t)=25*-1+31 -1 = smallest cosine h(t)=-25+31 h(t)=6ft

OpenStudy (johnweldon1993):

Exactly what I have as well...just took a longer path lol

OpenStudy (anonymous):

really? thank you !!!~ I have a question though, i dont know why everything that is in the parenthesis disapears

OpenStudy (johnweldon1993):

well see thats were my step would have come into play when you take the derivative and set it equal to 0...you get t = 14 and then when you plug in 14 into h(t) you get \[\large h(t) = 25cos(\frac{14\pi}{14}) + 31\] Which is \[\large h(t) = 25cos(\pi) + 31\] And we know cos(pi) = -1 so that gives us our \[\large h(t) = -25 + 31 = 6\]

OpenStudy (johnweldon1993):

Make sense now? :)

OpenStudy (anonymous):

totally makes sense now!!~ thank you so so so much!!~ would you mind helping me out on number 4 as well?

OpenStudy (johnweldon1993):

Yeah no problem :D This ones so easy all I'm gonna say is remember \[\large cos^2(\theta) = (1 - sin^2(\theta))\] And you should be all set :D but let me know if you're not!

OpenStudy (anonymous):

this is what I got for number 4 sinθ-sinθ*cos^2θ=sin^3θ 0*cos^2θ=sin^3θ 0=sin^3θ

OpenStudy (anonymous):

was this even right at all? lol

OpenStudy (johnweldon1993):

Nope lol...just look at it again quick its sin - (sin*cos^2) cant just subtract the sine's

OpenStudy (anonymous):

ok gimme a second to make an attempt to solve this lol

OpenStudy (anonymous):

ok idk how to multiply sinθ * cos^2θ in the parenthesis

OpenStudy (anonymous):

is that equal to something?

OpenStudy (anonymous):

because i know that there are 2 different identities

OpenStudy (anonymous):

but i dont know if they apply here or not :/

OpenStudy (johnweldon1993):

Ignore the parenthesis...I put them there purely to say dont just subtract... look a couple posts up about what I said about cos^2

OpenStudy (anonymous):

ohhh so i plug it in like this? sinθ-sinθ * (1-sin^2(θ))=sin^3θ

OpenStudy (johnweldon1993):

Perfect :) now just distribute the sin into the parenthesis and simplify

OpenStudy (anonymous):

do i convert them into 2 binomials to multiply them for distribution?

OpenStudy (johnweldon1993):

Just simple distribution \[\large sin - sin(1 - sin^2) = sin^3\] since -sin times 1 = -sin and -sin times -sin^2 = sin^3 this would become \[\large sin - sin + sin^3 = sin^3\]

OpenStudy (anonymous):

so would it take me here? sinθ-sin+sin^3=sin^3θ 0+sin^3=sin^3θ sin^3θ=sin^3θ

OpenStudy (anonymous):

i think im not getting it. im still confused

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

Okay so let me run through it (whole) we have \[\large sin(\theta) - sin(\theta)cos^2(\theta) = sin^3(\theta)\] The important identity to remember is \[\large sin^2 + cos^2 = 1\] we can rewrite that as \[\large cos^2 = 1 - sin^2\] So...looking back to our equation...lets replace \(\large cos^2\) with \(\large 1 - sin^2\) \[\large sin(\theta) - sin(\theta)cos^2(\theta) = sin^3(\theta)\] becomes \[\large sin(\theta) - sin(\theta)(1 - sin^2(\theta)) = sin^3(\theta)\] Now we just need to multiply \(\large sin(\theta)\) into the parenthesis, so multiply it by each term inside the parenthesis \[\large sin(\theta) - sin(\theta) + sin^3(\theta) = sin^3(\theta)\] And obviously we see that sin - sin = 0 so we are left with the confirmation \[\large sin^3(\theta) = sin^3(\theta)\]

OpenStudy (anonymous):

IM CONFUSED ABOUT WHAT YOU DID HERE IN THIS STEP: becomes sin(θ)−sin(θ)(1−sin2(θ))=sin3(θ) Now we just need to multiply sin(θ) into the parenthesis, so multiply it by each term inside the parenthesis sin(θ)−sin(θ)+sin3(θ)=sin3(θ)

OpenStudy (anonymous):

when you multiply them int he parenthesis, well this is what I got when I distribute them: sinθ-sin^3θ-sinθ+sin^3=sin^3θ after i simplify this, it also gives me 0=sin^3θ

OpenStudy (anonymous):

wait i just saw the - sign between them

OpenStudy (johnweldon1993):

Why do you get that? Oh I see what you did...no you're thinking again that there are parenthesis \[\large sin(\theta) - sin(\theta)(1 - sin^2(\theta))\] This is NOT the same as \[\large (sin(\theta) - sin(\theta))(1 - sin^2(\theta))\]

OpenStudy (anonymous):

sinθ-sinθ(1-sin^2θ)=sin^3θ sinθ-sinθ+sin^3θ=sin^3θ 0+sin^3θ=sin^3θ sin^3θ=sin^3θ

OpenStudy (anonymous):

this is it right? hahaha

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

That is perfect lol

OpenStudy (anonymous):

thank you :)

OpenStudy (johnweldon1993):

Great job! :)